OpenStudy (anonymous):

how would i compute the derivative f'(x), for f(x)=x 2/3

OpenStudy (lalaly):

\[\large{f(x)=x^{\frac{2}{3}}}\] is this ur question??

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

\[\huge{f'(x)=\frac{2}{3}x^{\frac{2}{3}-1}}\] use the rule\[\huge{f(x)=x^n}\]\[\huge{f'(x)=nx^{n-1}}\]

OpenStudy (anonymous):

\[2\large/3x^\frac{1}{3}\]